7. Network security software tracks the number of countries where software scams originate each day. Results from 2017 show that 2% of the legitimate software offers originate from two or more countries in a single day. In contrast, 28% of scam software originates software offers from two or more countries in a single day. Among all online software offers, the proportion of software scams is 0.02%. If the same software offer originates from two or more countries in a single day, what is the probability that the offer is a scam?
Answer #1

P(originate from 2 or more countries)=P(scam and originate from 2 or more countries)+P(not scam and originate from 2 or more countries)=0.0002*0.28+(1-0.0002)*0.02=0.020052

hence P(scam given originate from 2 or more countries)=P(scam and originate from 2 or more countries)/P(originate from 2 or more countries) =0.0002*0.28/0.020052=0.002793